# Forms of Understanding for XAI-Explanations

### v2

@@ 5.2 Influences brought by society
So far, the question of the potential impact of social processes on explanations and related processes of understanding has been excluded. However, how explainable AI – or any other new technology – is reflected upon in society has implications for the understanding of the individuals interacting with it, as humans constitute social reality by interpreting and creating their world and its objects in interactions [83]. Since the purpose of co-constructing Explainable Artificial Intelligence (XAI) is to support both machines and humans in their communication with each other, the understanding capabilities of both should be considered. The considerations relate to social constructionist approaches, according to which successful social acceptance, adoption and understanding of newly introduced technologies are highly dependent on the meaning co-constructed in social interactions, i.e., when human (or non-human) agents or group members interact. In this context, group membership (e.g., constituted by gender, class, race, age, socioeconomic status) plays a key role in the process of understanding technological novelty and thus in explaining it. Group membership “determines what is relevant and what is not” [60, p. 361]. In other words, goals or relationships may be different for different social groups.
The social constructivist view of human-human interaction is also relevant to human-machine interaction, and even precedes it. Ehsan and Riedl [33] emphasize the importance of a human-centered approach to XAI. Following their ‘reflective sociotechnical approach’, we propose a concept of reflective understanding that critically reflects the dominant narratives as well as the explicit and implicit assumptions and practices of the field. Moreover, it is value-sensitive to researchers, designers, and users of AI, including to those who tend to be excluded from the conversation and are denied access to the technology [33, p. 461]. In other words, in order to understand how and why people act the way they do (e.g., adoption or non-adoption of a particular new technology) social meanings, interactions and interpretations cannot be ignored. Accordingly, explanations must also address how people interpret their world and how they interactively create this world. Not only the perspectives and stocks of knowledge that we are aware of should be of interest when designing an explanation, but also implicit knowledge and the interactive creation of meanings that people (users) are usually not aware of [83].
@@
Applying our concept of agency (see Section 3.4), a high degree of comprehension and enabledness is characterized by healthy distrust [90], deep enablement, knowledge of the strengths and power, but also of the limitations, biases, and potential for and actualization of oppression and exclusion that AI technology entails. In addition, agency means being empowered to overcome socially and technologically consolidated marginalization/discrimination that may be related to group membership. Thus, in order to support agency, explainers (human or XAI) should take into account explainees’ group-related inequalities and differences, as well as varying access to resources and positions/roles in socio-technical systems, because group affiliation “determines what is relevant and what is not” [60, p. 361]. Indeed, we see this as inevitable for AI development because a conceptualization of human understanding of XAI-explanations without context is insufficient, and concepts of human understanding should be central to the research and design of XAI methods and technology so that they can do justice to the humans they aim to empower.

### v3

@@ 6.2 Influences brought by society
So far, the question of the potential impact of social processes on explanations and related processes of understanding has been excluded. However, how explainable AI – or any other new technology – is reflected upon in society has implications for the understanding of the individuals interacting with it, as humans constitute social reality by interpreting and creating their world and its objects in interactions (Rosenthal, 2018). Since the purpose of co-constructing Explainable Artificial Intelligence (XAI) is to support both machines and humans in their communication with each other, the understanding capabilities of both should be considered. The considerations relate to social constructionist approaches, according to which successful social acceptance, adoption and understanding of newly introduced technologies are highly dependent on the meaning co-constructed in social interactions, i.e., when human (or non-human) agents or group members interact. In this context, group membership (e.g., constituted by gender, class, race, age, socioeconomic status) plays a key role in the process of understanding technological novelty and thus in explaining it. Group membership “determines what is relevant and what is not” (Kronberger, 2015, p. 361). In other words, goals or relationships may be different for different social groups.
The social constructivist view of human interaction is also relevant to human-machine interaction, and even precedes it. Ehsan and Riedl (2020) emphasize the importance of a human-centered approach to XAI. Following their ‘reflective sociotechnical approach’, we propose a concept of reflective understanding that critically reflects the dominant narratives as well as the explicit and implicit assumptions and practices of the field. Moreover, it is value-sensitive to researchers, designers, and users of AI, including to those who tend to be excluded from the conversation and are denied access to the technology (Ehsan and Riedl, 2020, p. 461). In other words, in order to understand how and why people act the way they do (e.g., adoption or non-adoption of a particular new technology) social meanings, interactions and interpretations cannot be ignored. Accordingly, explanations must also address how people interpret their world and how they interactively create this world. Not only the perspectives and stocks of knowledge that we are aware of should be of interest when designing an explanation, but also implicit knowledge and the interactive creation of meanings that people (users) are usually not aware of (Rosenthal, 2018).
@@
Applying our concept of agency (see Section 3.4), a high degree of comprehension and enabledness is characterized by healthy distrust (Visser et al., 2025), deep enablement, knowledge of the strengths and power, but also of the limitations, biases, and potential for and actualization of oppression and exclusion that AI technology entails. In addition, agency means being empowered to overcome socially and technologically consolidated marginalization/discrimination that may be related to group membership. Thus, in order to support agency, explainers (human or XAI) should take into account explainees’ group-related inequalities and differences, as well as varying access to resources and positions/roles in socio-technical systems, because group affiliation “determines what is relevant and what is not” (Kronberger, 2015, p. 361). Indeed, we see this as inevitable for AI development because a conceptualization of human understanding of XAI-explanations without context is insufficient, and concepts of human understanding should be central to the research and design of XAI methods and technology so that they can do justice to the humans they aim to empower.